OpenStudy (anonymous):

(x+8)(x+8) solution?

OpenStudy (anonymous):

x^2+2*x*8+8*8 x2+16x+64

OpenStudy (anonymous):

idk if its 8,0 or -8,0 @khurramshahzad

OpenStudy (anonymous):

wait have some work get u after 30 minuts

OpenStudy (texaschic101):

are you familiar with the foil method ?

OpenStudy (texaschic101):

(x + 8)(x + 8) FOIL (first,outside,inside,last) F - multiply the first terms in each set x * x = x^2 O - multiply outside terms in each set x * 8 = 8x I - multiply inside terms in each set 8 * x = 8x L - multiply the last terms in each set 8 * 8 = 64 put it all together x^2 + 8x + 8x + 64 combine like terms x^2 + 16x + 64......here you go Do you have any questions ?